OpenStudy (anonymous):

China has a population of approximately 1.3 x 10^9 people. The United states has a population of approximately 3 x 10^8 people. How many more people live in China than in the United States?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

\[\frac{1.3\times 10^9}{3\times 10^8}=\frac{1.3}{3}\times 10^{9-8}=\frac{1.3}{3}\times 10\]

OpenStudy (anonymous):

k . thanks !